# Post-Newtonian dynamics of compact binaries with mass transfer

### v1

@@ I Introduction
This paper is organized as follows. In Section II, we present the post-Newtonian retarded potential $V$ with MT in the metric at the 2PN order. In Section III, we derive the Fokker action with MT at the 2PN order. In Section IV, we obtain the relative acceleration of compact binaries at the 2.5PN order. In Section V, we derive the variation of the frequency of GWs in the case of quasi-circular orbits and analyze the contribution of the MT effect at different PN orders. Throughout, we use the Minkowski metric $\eta_{\mu\nu}$ with signature (,+,+,+)(-,+,+,+)( - , + , + , + ) and choose the harmonic gauge $\partial_{\mu}(g^{\mu\nu}\sqrt{-g})=0$.
@@
We will derive the specific expressions of the metric $g_{\mu\nu}$, the energy-momentum tensor $T^{\mu\nu}$, and the PN potential $V_{\mu}=(V,V_{i})$ with the MT effect at the 2PN order. Considering MT, the 2PN metric takes the form as blanchet2024postnewtoniantheorygravitationalwaves
@@
which is the same as the classical PN theory without MT Blanchet2014 ; PhysRevD.63.062005 . This PN metric is expanded in powers of $1/c$, with $c$ the speed of light. The PN potentials $V,W,R,X$ are determined by PhysRevD.63.062005 ; PhysRevD.81.064004
@@ III Action and Lagrangian
where $S_{g}$ is the gravitational action written in the Landau and Lifshitz form landau2013classical , $S_{m}$ is the matter action takes from the N-body system, and $R$ is the Ricci Scalar. $\Gamma^{\alpha}\equiv g^{\mu\nu}\Gamma^{\alpha}_{\mu\nu}$ and $-\frac{1}{2}\sqrt{-g}g_{\mu\nu}\Gamma^{\mu}\Gamma^{\nu}$ is the standard gauge-fixing term appropriate to harmonic coordinates. $v_{A}^{\mu}\equiv\mathrm{d}x_{A}^{\mu}/\mathrm{d}t=(c,\bm{v}_{A})$ is the coordinate velocity of particle $A$ and $(g_{\mu\nu})_{A}$ is the metric evaluated at the location of the particle $A$. $S_{g}$ can be written as
@@
$\begin{split}c^{4}\sqrt{-g}J=&-2V\nabla^{2}V+\frac{1}{c^{2}}\left\{2V\ddot{V}+8V^{i}\nabla^{2}V_{i}\right\}\\ &+\frac{1}{c^{4}}\left\{-12V^{2}\ddot{V}+8W^{i}_{i}\ddot{V}+32R_{i}\nabla^{2}V^{i}+16W^{i}_{i}\ddot{V}-16\dot{V}^{i}\partial^{k}W_{ik}+16VV_{i}\nabla^{2}V^{i}+4W^{i}_{i}V_{,k}V^{,k}\right.\\ &\qquad\quad-8W^{j}_{i}V^{,i}V_{,j}+32R_{i}\nabla^{2}V^{i}+32R^{k}_{,k}V^{i}_{,i}-32V_{i}V_{j}V^{,i,j}-32V_{j}V^{i}_{,i}V_{,j}-16X\nabla^{2}V\\ &\qquad\quad\left.-8W^{j}_{i}\nabla^{2}W^{j}_{i}+8W^{i}_{i}\nabla^{2}W^{j}_{j}-8W^{k}_{k}W_{ij}^{,i,j}-8W_{j}^{i,j}W_{ik}^{,k}\right\}+\mathcal{O}(c^{-6})\\ &+\text{the terms of total derivative}\left[\partial_{t}(...),\partial_{i}(...)\right],\end{split}$ (23)
$\begin{split}c^{4}\frac{1}{2}\sqrt{-g}g_{\mu\nu}\Gamma^{\mu}\Gamma^{\nu}=&\frac{1}{c^{4}}\left\{2V_{i}\nabla^{2}V^{i}+2V_{i,j}V^{i,j}+16\dot{V}^{i}W_{i,j}^{j}-8\dot{V}^{i}W_{j,i}^{j}-8W_{i,j}^{j}W_{,k}^{ik}+8W_{i,j}^{j}W_{k,i}^{k}-2W_{j,i}^{j}W_{k,i}^{k}\right\}\\ &+\mathcal{O}(c^{-6})+\text{the terms of total derivative}\left[\partial_{t}(...),\partial_{i}(...)\right].\end{split}$ (24)
Here $(...)_{,\mu}\equiv\partial_{\mu}(...)$ and $\partial_{0}=c^{-1}\partial_{t}$. We have used the relation $V^{i}_{,i}+cV_{,0}=0$ which is due to the continuity equation. The contribution of terms of gauge-fixing start from the 2PN order. From Eq.(A) we find that the effect caused by the variation in mass come from the order of $c^{-2}$ in the PN potential $V$. Terms with the variation in mass at the 2PN order can be calculated by the 1PN potential. Other 2PN potentials can be calculated the as classical PN method since their effects incorporate the variation in mass start from the 3PN order.
@@
the terms without MT is same as the Lagrangian in Eq.(4.1) in VanessaCdeAndrade_2001 . In non-extreme cases (excluding violent accretion or extreme eccentricity), the MT rate is typically considered to be small. At the frequency of 10 mHz and with Eddington accretion, the contribution of MT at the Newtonian-order is comparable to that of GW radiation PhysRevD.111.043049 . Therefore, terms of the order $\dot{m}^{2}$ can be regarded as a higher-order quantity and be neglected in the calculation. Additionally, in cases without extreme eccentricity, the variation in the MT rate is not significant, implying that the second derivative of masses $\ddot{m}$ can also be treated as a small quantity and neglected.
@@
This form is much more simpler than Eq.(26). We will use this Lagrangian to derive angular momentum and angular acceleration in the next section. One peculiarity is that all MT terms include the dimensionless relative mass difference [^1] Note that the symbols for the relative mass difference $\Delta$ and the Laplace operator $\bigtriangleup$ are different. $\Delta$ and the MT rate $\chi$
@@ IV.1 The 2PN acceleration in harmonic-coordinate
[…]
@@
[…]
@@ IV.3 Center-of-mass frame
~$\tilde{\mathcal{B}}_{0}$ $=\frac{1}{c^{2}}\bigg{\{}-4\dot{r}+2\dot{r}\eta\bigg{\}}+\frac{1}{c^{4}}\bigg{\{}\frac{9}{2}\dot{r}^{3}\eta+3\dot{r}^{3}\eta^{2}-\frac{15}{2}\dot{r}\eta v^{2}-2\dot{r}\eta^{2}v^{2}+\frac{GM}{r}\left(2\dot{r}+\frac{41}{2}\dot{r}\eta+4\dot{r}\eta^{2}\right)\bigg{\}}$
@@
~$\tilde{\mathcal{B}}_{MT}$ $=\frac{1}{c^{2}}\bigg{\{}-1\bigg{\}}+\frac{1}{c^{4}}\bigg{\{}-3\eta^{2}\dot{r}^{2}-6\eta\dot{r}^{2}+\frac{3}{4}\dot{r}^{2}-3\eta^{2}v^{2}+6\eta v^{2}+\frac{13}{4}v^{2}+\frac{GM}{r}\left(-\eta^{2}+\frac{29}{4}\eta+\frac{3}{4}\right)\bigg{\}}$
@@ V Evolution of the frequency in circular orbits
$+\frac{1}{c^{5}}\bigg{\{}1+\frac{1}{c^{2}}(GM)^{2/3}\bigg{(}-\frac{743}{336}-\frac{11}{4}\eta\bigg{)}\omega^{2/3}\bigg{\}}\frac{96}{5}(GM)^{5/3}\omega^{8/3}.$ (54)
@@
the MT rate $\chi$ $1\times 10^{-5}M_{\odot}/\text{yr}$ $1\times 10^{-6}M_{\odot}/\text{yr}$
Frequency $f_{0}\;(\text{mHz})$ $100.0\thicksim 163.0$ $10\thicksim 10-1.2\times 10^{-8}$ $100.0-163.2$ $10\thicksim 10+5\times 10^{-6}$
@@
Total - Newtonian $-1.56\times 10^{3}$ $-0.087$ $-1.56\times 10^{3}$ $-0.045$
Newtonian MT $-5.25\times 10^{3}$ $-4.66\times 10^{2}$525.13 $-46.56$
1PN MT $-2.68$ $-0.047$0.268 $-4.70\times 10^{-3}$
2PN MT      $\;\;\;\;\;2.41\times 10^{-3}$      $\;\;\;\;\;8.32\times 10^{-6}$      $\;\;\;\;\;2.41\times 10^{-4}$ $-8.32\times 10^{-7}$
@@
1PN GW $-1.56\times 10^{3}$ $-0.040$ $-1.56\times 10^{3}$ $-0.040$
In Table 1, by relying on Eq.(V), numerical integration is employed to compute the contribution of each term to the number of cycles of GWs within quasi-circular binary systems after a five-year evolutionary period. The “Total - Newtonian” represents the total effect minus the Newtonian effect, which have already been computed in PhysRevD.111.043049 . At the frequency of 100 mHz, the effect of MT at the 1PN order reaches $-2.68$ cycles, which have a detectable signal on GW detection due to the long-term cumulative results. At the frequency 10 mHz, the contributions of MT and GW radiation are comparable.
@@ Appendix C Transformation to the center-of-mass frame
$E_{A}=$ $m_{A}c^{2}+\left\{\frac{1}{2}m_{A}v_{A}^{2}-\frac{Gm_{A}m_{B}}{2r}\right\}+\frac{1}{c^{2}}\left\{\frac{3}{8}m_{A}v_{A}^{4}+\frac{Gm_{A}m_{B}}{r}\left(\frac{19}{8}v_{A}^{2}-\frac{7}{8}v_{B}^{2}-\frac{5}{4}\frac{Gm_{A}}{r}+\frac{7}{4}\frac{Gm_{B}}{r}\right.\right.$

### v2

@@ I Introduction
This paper is organized as follows. In Section II, we present the post-Newtonian retarded potential $V$ with MT in the metric at the 2PN order. In Section III, we derive the Fokker action with MT at the 2PN order. In Section IV, we obtain the relative acceleration of compact binaries at the 2.5PN order. In Section V, we derive the variation of the frequency of GWs in the case of quasi-circular orbits and analyze the contribution of the MT effect at different PN orders. Throughout, we use the Minkowski metric $\eta_{\mu\nu}$ with signature (,+,+,+)(-,+,+,+) and choose the harmonic gauge $\partial_{\mu}(g^{\mu\nu}\sqrt{-g})=0$.
@@
We will derive the specific expressions of the metric $g_{\mu\nu}$, the energy-momentum tensor $T^{\mu\nu}$, and the PN potential $V_{\mu}=(V,V_{i})$ with the MT effect at the 2PN order. Considering the MT, the 2PN metric takes the form as blanchet2024postnewtoniantheorygravitationalwaves
@@
which is also applicable to the binary system with MT Gravity_PoissonWill ; Blanchet2014 ; PhysRevD.63.062005 . This PN metric is expanded in powers of $1/c$, with $c$ the speed of light. The PN potentials $V,W,R,X$ are determined by PhysRevD.63.062005 ; PhysRevD.81.064004
@@ III Action and Lagrangian
where $S_{g}$ is the gravitational action written in the Landau and Lifshitz form landau2013classical , $S_{m}$ is the matter action takes from the N-body system, and $R$ is the Ricci Scalar. $\Gamma^{\alpha}\equiv g^{\mu\nu}\Gamma^{\alpha}_{\mu\nu}$ and $\frac{1}{2}\sqrt{-g}g_{\mu\nu}\Gamma^{\mu}\Gamma^{\nu}$ is the standard gauge-fixing term appropriate to harmonic coordinates. $v_{A}^{\mu}\equiv\mathrm{d}x_{A}^{\mu}/\mathrm{d}t=(c,\bm{v}_{A})$ is the coordinate velocity of particle $A$ and $(g_{\mu\nu})_{A}$ is the metric evaluated at the location of the particle $A$. $S_{g}$ can be written as
@@
[…]
@@
Through our derivation, we found that the Lagrangian $\mathcal{L}_{0}$ without MT is consistent with the Lagrangian in Eq.(4.1) in VanessaCdeAndrade_2001 . Meanwhile, additional effects caused by MT in Lagrangian $\mathcal{L}_{MT}$ are obtained. In non-extreme cases (excluding violent accretion or extreme eccentricity), the MT rate is typically considered to be small. At a frequency of 10 mHz and a MT rate of $10^{-5}M_{\odot}/\text{year}$, the contribution of $\dot{m}$ at the Newtonian-order is comparable to that of GW radiation at the 2.5 PN order PhysRevD.111.043049 and the contribution of $\dot{m}^{2}$ is much smaller than that of GW at the 2.5 PN order. Therefore, terms of the order $\dot{m}^{2}$ can be regarded as a higher-order quantity and be neglected in the calculation. Additionally, in cases without extreme eccentricity, the variation in the MT rate is not significant, implying that the second derivative of masses $\ddot{m}$ can also be treated as a small quantity and neglected.
@@
This form is much more simpler than Eq.(26). We will use this Lagrangian to derive angular momentum and angular acceleration in the next section. One peculiarity is that all MT terms include the dimensionless relative mass difference [^1] Note that the symbols for the relative mass difference $\Delta$ and the Laplace operator $\bigtriangleup$ are different. $\Delta$ and the MT rate $\chi$.
@@ IV.1 The 2PN acceleration in harmonic-coordinate
$a^{i}_{2.5PN}=\frac{G}{c^{5}}\left[-\frac{GM}{r^{2}}\left(3{\mathop{I}^{(3)}}_{jk}n^{j}n^{k}+\frac{1}{3}\mathop{I}^{(3)}{}^{k}_{k}\right)n^{i}+2\mathop{I}^{(4)}{}^{i}_{k}v^{k}+\frac{3}{5}\left(\mathop{I}^{(5)}{}^{i}_{k}-\frac{1}{3}\delta^{i}_{k}\mathop{I}^{(5)}{}^{j}_{j}\right)r^{k}\right],$ (37)
@@
$\mathop{I}^{(3)}{}^{jk}_{MT}$= $\mathop{I}^{(3)}{}^{jk}_{0}-4\frac{GM}{r}\dot{\mu}n^{j}n^{k}+\mathcal{O}(\dot{\mu}^{2}),$ (40)
$\mathop{I}^{(4)}{}^{jk}_{MT}$= $\mathop{I}^{(4)}{}^{jk}_{0}-\frac{GM}{r^{2}}\bigg\{4\bigg[(n^{j}v^{k}+n^{k}v^{j})-3\dot{r}n^{k}n^{j}\bigg]\dot{\mu}\bigg\}+\mathcal{O}(\dot{\mu}^{2}),$ (41)
$\begin{split}\mathop{I}^{(5)}{}^{jk}_{MT}=&\mathop{I}^{(5)}{}^{jk}_{0}+2\frac{GM}{r^{3}}\left\{\left[\frac{GM}{r}n^{j}n^{k}-3\left(v^{2}n^{j}n^{k}+\dot{r}(n^{j}v^{k}+n^{k}v^{j})-5\dot{r}^{2}n^{j}n^{k}\right)\right]\dot{\mu}\right\}+\mathcal{O}(\dot{\mu}^{2}),\end{split}$ (42)
@@ IV.3 Center-of-mass frame
$\tilde{\mathcal{B}}_{0}$ $=\frac{1}{c^{2}}\bigg\{-4\dot{r}+2\dot{r}\eta\bigg\}+\frac{1}{c^{4}}\bigg\{\frac{9}{2}\dot{r}^{3}\eta+3\dot{r}^{3}\eta^{2}-\frac{15}{2}\dot{r}\eta v^{2}-2\dot{r}\eta^{2}v^{2}+\frac{GM}{r}\left(2\dot{r}+\frac{41}{2}\dot{r}\eta+4\dot{r}\eta^{2}\right)\bigg\}$
@@
$\tilde{\mathcal{B}}_{MT}$ $=\frac{1}{c^{2}}\bigg\{-1\bigg\}+\frac{1}{c^{4}}\bigg\{-3\eta^{2}\dot{r}^{2}-6\eta\dot{r}^{2}+\frac{3}{4}\dot{r}^{2}-3\eta^{2}v^{2}+6\eta v^{2}+\frac{13}{4}v^{2}+\frac{GM}{r}\left(-\eta^{2}+\frac{29}{4}\eta+\frac{3}{4}\right)\bigg\}$
@@ V Evolution of the frequency in circular orbits
$+\frac{1}{c^{5}}\bigg\{1+\frac{1}{c^{2}}(GM)^{2/3}\bigg(-\frac{743}{336}-\frac{11}{4}\eta\bigg)\omega^{2/3}\bigg\}\frac{96}{5}(GM)^{5/3}{\eta}\omega^{8/3}.$ (54)
@@
the MT rate $\chi$ $\times 10^{-5}M_{\odot}/\text{yr} 1×1$$\times 10^{-6}M_{\odot}/\text{yr}$
Frequency $f_{0}\;(\text{mHz})$ $0.0\thicksim 163.0 1010$$\thicksim 10-1.2\times 10^{-8}$ 100.0163.2100.0-163.2 $10\thicksim 10+5\times 10^{-6}$
@@
Total - Newtonian $1.56\times 10^{3}$ 0.087-0.087 $1.56\times 10^{3}$ 0.045-0.045
Newtonian MT $5.25\times 10^{3}$ $4.66\times 10^{2} -$525.13 46.56-46.56
1PN MT 2.68-2.68 0.047-0.047 -0.268 $4.70\times 10^{-3}$
2PN MT      $2.41\times 10^{-3}$      $8.32\times 10^{-6}$      $2.41\times 10^{-4}$ $8.32\times 10^{-7}$
@@
1PN GW $1.56\times 10^{3}$ 0.040-0.040 $1.56\times 10^{3}$ 0.040-0.040
In Table 1, by relying on Eq.(V), numerical integration is employed to compute the contribution of each term to the number of cycles of GWs within quasi-circular binary systems after a five-year evolutionary period. The “Total - Newtonian” represents the total effect minus the Newtonian effect, which have already been computed in PhysRevD.111.043049 . At the frequency of 100 mHz, the effect of MT at the 1PN order reaches 2.68-2.68 cycles, which have a detectable signal on GW detection due to the long-term cumulative results. As shown in Fig. 1, for a typical binary system, the MT contribution at Newtonian order is much smaller than the GW contribution at Newtonian order, but larger than the GW contribution at 1PN order.
@@ Appendix C Transformation to the center-of-mass frame
[…]